OpenStudy (anonymous):

We have a wooden piece with a volumetric mass of 800kg/m^3 and a mass of 200c.m^2 that floats on water. Find the submerged volume of the piece. (Knowing that the volumetric mass of water is 1000kg/m^3)

OpenStudy (anonymous):

Please word this question correctly. Your wording is conflicting.

OpenStudy (anonymous):

Well i don't know how to explain it because english is my second language

OpenStudy (anonymous):

I wish if I could post a picture

OpenStudy (anonymous):

The question is to find the volume of the submerged part of the wooden piece

OpenStudy (irishboy123):

"volumetric mass " = density what does this mean: "200c.m^2"

OpenStudy (irishboy123):

weight or mass? they are edifferent. do you mean 100g ?? 0.1kg?? c.m^2 does not mean anything obvious

OpenStudy (anonymous):

Soooooorrrryy my bad, it's volume i meant

OpenStudy (anonymous):

You have to change it into m^3

OpenStudy (irishboy123):

so its volume is 200. 200 what? you wrote c.m^2 for your units.

OpenStudy (anonymous):

Yea yes you have to change it into cubic meter

OpenStudy (irishboy123):

the relevant principle is : A floating object displaces its own mass of the fluid in which it floats. so for a body (b) in water (w) we can say \(m_b = m_w\) for the mass of body and the mass of displaced water thus \(\rho_b * V_b = \rho_w* V_w\) where \(\rho\) is density and \(V\) is volume plug in numbers

OpenStudy (anonymous):

Pb =800kg/m^3. Pw=1000kg/m^3. Vb= 200cm^3=200*10^-6m^3

OpenStudy (irishboy123):

\(\huge V_w = \frac{ρ_b∗V_b}{ρ_w} \)